OpenStudy (anonymous):

Simplify: (4xy^-2)/(12x^-1/3y^-5)

myininaya (myininaya):

\[\frac{4xy^{-2}}{\frac{12x^{-1}}{3y^{-5}}}\] is that right?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Sorry I don't know how to make it look like a real equation

myininaya (myininaya):

\[4xy^{-2} \div \frac{12x^{-1}}{3y^{-5}}=4xy^{-2} \times \frac{3y^{-5}}{12x^{-1}}\] \[\frac{4xy \cdot 3y^{-5}}{12x^{-1}}=\frac{12xy^{-5+1}}{12x^{-1}}=\frac{xy^{-4}}{x^{-1}}\] \[=\frac{x}{x^{-1}} \cdot \frac{y^{-4}}{1}\] \[=x \cdot x \frac{1}{y^4}=x^2 \frac{1}{y^4}=\frac{x^2}{y^4}\]

OpenStudy (anonymous):

Wonderful thank you so much

myininaya (myininaya):

np i tried to put in alot of steps so it was easy to follow

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (anonymous):

4xy^-2/12x^-1/3 y^-5